The perimeter of a rectangular flower garden is 24cm if the dimensions are quadrupled what would be the perimeter of the larger garden
Answer by Boreal(15235)

2L+2W=24
L+W=12
4L+4W=48
It is doubled.
10 and 2 give a perimeter of 24.
40 and 8 give a perimeter of 48 and the dimensions have been quadrupled.
(The area would be 16 times as much.)